Malek drove for 2 hours then stopped for a break. After that, he drove for 3 hours at a
speed 10 miles per hour (mph) more than the speed before the break. The total distance
he drove was 280 miles. What was Malek’s speed before the break?

Answer by stanbon(57328)

Malek drove for 2 hours then stopped for a break.
After that, he drove for 3 hours at a speed 10 miles per hour (mph) more than the speed before the break.
The total distance he drove was 280 miles.
What was Malek’s speed before the break?
---------------------------------------------
Before break:
time = 2 hrs ; rate = x mph ; distance = 2x miles
---
After break:
time = 3 hrs ; rate = x+10 mph ; distance = 3x+30 miles
-------------------------------
Equation:
distance + distance = 280
2x + 3x+30 = 280
5x = 250
x = 50 mph (rate before the break)
======================================
cheers,
Stan H.
==============